I need to solve for t given the rational equation, 1/t+5/(t+1)=10
Answer by stanbon(75887)

solve for t given the rational equation, 1/t+5/(t+1)=10
------------
Multiply thru by t(t+1) to get:
t+1 + 5t = 10t(t+1)
6t + 1 = 10t^2+10t
---
10t^2+4t-1 = 0
----
t = [-4 +- sqrt(100-4*10*-1)]/20
-----
t = [-4 +- sqrt(140)]/20
-----
t = -0.3916 or x = -0.7916
=============================
Cheers,
Stan H.
============